OpenStudy (anonymous):

4x+6+3=17 multi-step equations

jigglypuff314 (jigglypuff314):

to solve for x for: 4x+6+3=17 first simplify what you can so 4x + 9 = 17 then try to get x alone on one side so -9 -9 4x = 8 then divide both sides by 4 /4 /4 x = 8/4 which can be simplified to get your answer :)